OpenStudy (anonymous):

Only two forces act on an object. Find the magnitudeand direction (relative to the x axis) of the acceleration of the object. Mass=3.0kg F1=40.0 N due east F2=60 N, 45 degrees North of east The answer is: 30.9 m/s 2, 27.2 deg above the x axis How do I get this answer?

OpenStudy (anonymous):

60(cos45) and 60(sin45) are both 42.4 Total X is 82.4 Total Y is 42.4 Is this correct so far?

OpenStudy (anonymous):

yes it is :) I'm surprised that no one has confirmed that.

OpenStudy (anonymous):

what do I do next?

OpenStudy (anonymous):

Net force=\(\sqrt(F_x^{2}+F_y^{2})\)

OpenStudy (anonymous):

\[\large tan\theta=\frac{F_y}{F_x}\]